# DECODE: Domain-aware Continual Domain Expansion for Motion Prediction

### v1

@@ Abstract
Motion prediction is critical for autonomous vehicles to effectively navigate complex environments and accurately anticipate the behaviors of other traffic participants. As autonomous driving continues to evolve, the need to assimilate new and varied driving scenarios necessitates frequent model updates through retraining. To address these demands, we introduce DECODE, a novel continual learning framework that begins with a pre-trained generalized model and incrementally develops specialized models for distinct domains. Unlike existing continual learning approaches that attempt to develop a unified model capable of generalizing across diverse scenarios, DECODE uniquely balances specialization with generalization, dynamically adjusting to real-time demands. The proposed framework leverages a hypernetwork to generate model parameters, significantly reducing storage requirements, and incorporates a normalizing flow mechanism for real-time model selection based on likelihood estimation. Furthermore, DECODE merges outputs from the most relevant specialized and generalized models using deep Bayesian uncertainty estimation techniques. This integration ensures optimal performance in familiar conditions while maintaining robustness in unfamiliar scenarios. Extensive evaluations confirm the effectiveness of the framework, achieving a notably low forgetting rate of 0.044 and an average minADE of 0.584 m, significantly surpassing traditional learning strategies and demonstrating adaptability across a wide range of driving conditions.
Index Terms:
Motion Prediction, continual learning, domain awareness, autonomous driving.
@@ I Introduction
Motion prediction has attracted significant attention in recent years, finding applications across a wide array of fields. It plays a critical role in the development of autonomous vehicles, enabling them to understand complex scenarios and anticipate the future behaviors of other traffic participants [1, 2, 3, 4, 5]. This capability allows for more informed decision-making, which is crucial not only for enhancing roadway safety but also for improving traffic efficiency and reducing congestion. Beyond road transport, motion prediction is instrumental in areas such as robotics for navigation and interaction [6, 7], and in simulation environments where predictive models support the development and testing of various autonomous systems [8, 9]. The task of motion prediction is inherently challenging due to the complexity of interactions among target agents and the multi-modal nature of their possible future movements.
@@
Recent advancements in state-of-the-art algorithms have focused on integrating complex multi-agent interactions and map data [10, 11, 12, 13, 14, 15], yielding successful performances on several prominent open datasets [16, 17, 18]. However, these models require extensive datasets covering a broad spectrum of interactions involving multiple agents and road topology. Recent studies [19] have shown that without further training, models that excel on one dataset may perform poorly on others, posing a challenge for autonomous vehicles’ global deployment and operation in diverse, unforeseen scenarios. Traditionally, motion prediction development has involved training models on entire datasets—a method still prevalent in current state-of-the-art efforts. However, the nature of autonomous driving development, characterized by the continuous acquisition of new and varied driving scenarios, necessitates frequent updates through model retraining. Retraining on all data combined becomes progressively more resource-intensive, while retraining solely on new data can lead to the loss of previously acquired knowledge [20]. This situation underscores the critical need for ongoing innovation in motion prediction methodologies and the development of effective continual learning strategies to enhance vehicle safety and reliability across varied environments. Therefore, in this paper we propose a novel continual learning framework for motion prediction called Domain-awarE COntinual Domain Expansion (DECODE).
@@
Extensive experiments confirm the effectiveness and superior performance of the DECODE framework. By leveraging a hypernetwork for continual learning, we achieved a notably low forgetting rate of 0.044, surpassing several other active learning strategies, and an average minADE of just 0.584 meters. Our normalizing flow-based model identification method accurately distinguishes among different domains, providing robust evidence of its ability to align model selection with specific operational scenarios, achieving an AUROC of 0.988. We further validated our framework using both naturalistic trajectory data and synthetically generated data, demonstrating its capacity to consistently produce high-quality motion predictions across a diverse range of domains, including safety-critical scenarios, with an accuracy of 0.986.
@@ II-A Motion Prediction
Neural network-based learning approaches have significantly outperformed traditional model-based methods such as constant velocity models and Kalman filters in motion prediction [29]. These neural networks typically employ an encoder-decoder structure. During the encoding phase, techniques such as recurrent neural networks and 1D convolutions take advantage of the sequential nature of data to encode past trajectories. Recent research has further advanced motion prediction by focusing on two critical scene contexts: multi-agent interactions and local map features. Initial pioneering efforts, including Social LSTM [30] and Social GANs [31], explored the impact of multi-agent interactions on future behaviors. Meanwhile, local maps have typically been rasterized into 2D images [32, 33, 34, 35, 36, 37, 38, 39, 10, 1], allowing CNNs to extract detailed scene representations from them.
@@
Subsequent advances in Graph Neural Networks (GNNs) and transformers have refined the capture of scene context information. LaneGCN [11] effectively discretizes the local map into lane nodes along lane centerlines, creating a connected lane graph that utilizes graph convolutions to extract relational information among agents and lanes—an approach echoed by several later models [40, 14]. Additionally, VectorNet [41], widely adopted in recent research [12, 13, 42, 43], encodes both agents and map elements as polylines into a global graph. A graph neural network is then employed to encode this graph, effectively integrating the spatial relationships within the scene. Originally utilized in natural language processing, transformers have recently gained significant traction in motion prediction [44, 45, 46, 47, 3, 15]. Like graph neural networks, they can flexibly extract relational information among agents and map elements by leveraging the multi-head attention mechanism [48].
@@
Given the possibility of multiple future outcomes, many studies have incorporated multi-modal predictions to approximate the distribution of potential future motions of a target agent. These approaches typically generate a discrete set of future predictions, each with a corresponding confidence score to capture outcome variability. Early methods [33, 34, 11, 49] used multiple regression heads to produce unique modes. Generative models like GANs [31, 50, 51] and CVAEs[52, 10, 2] are also widely employed. More recent methods [53, 54, 55] often involve an intermediate classification phase to identify a distinct future mode, followed by regression to refine the motion. For example, TNT [12] and denseTNT [13] use agent’s target endpoints to represent future modes, GOHOME [14] predicts future motions from a heatmap, and MTR++ [15] employ a learnable intention query with k-means clustered intention points. Our DECODE framework is compatible with any motion prediction model that utilizes an encoder-decoder structure and is flexible regarding input format, allowing for the integration of state-of-the-art scene context encoding methods. Furthermore, our framework adeptly handles the mult-modality of predictions by acknowledging the potential for multiple viable future trajectories and is particularly effective with methods that incorporate an intermediate classification phase.
@@ II-B Continual Learning
Continual learning can be broadly categorized into four approaches. Parameter isolation methods, such as progressive neural networks [56], expand untouched model parameters with the arrival of new domain data, completely protecting previous knowledge at the cost of increased model complexity and parameter storage. Regulation-based methods mitigate forgetting by adding a regularization term to the loss function; an example is Elastic Weight Consolidation (EWC) [57], which employs Fisher information to protect important parameters related to old knowledge during training. Knowledge distillation-based methods, such as Learning without Forgetting (LwF) [58], leverage outputs from old models to distill knowledge into the new model, preserving prior learning while integrating new information. Lastly, rehearsal-based methods include experience replay, which utilizes a selective collection from older datasets, and generative replay, which uses synthetically generated samples from previously learned models.
@@
These existing approaches present several drawbacks. Experience replay struggles with selecting representative samples and necessitates external storage, which can quickly become unmanageable [61]. Generative replay’s effectiveness heavily relies on the quality of synthetic data, posing significant challenges for state-of-the-art motion prediction models that require handling complex inputs like varying agent numbers and intricate road geometry. Moreover, training one AI model using synthetically generated data from another can degrade performance, a risk highlighted by Shumailov et al. in a recent Nature study [62]. To address these issues, our DECODE framework employs hypernetworks [25], which innovatively combine parameter isolation with the regularization of loss function. This approach efficiently preserves the integrity of previous models’ parameters while integrating new knowledge, offering a scalable and effective solution to the challenges of continual learning in motion prediction.
@@ III-A Hypernetworks for Continual Domain Expansion
$L_{\texttt{reg}}=\lambda\sum_{i=1}^{m-1}||\mathcal{H}(q^{(i)};\Theta^{*})-\mathcal{H}(q^{(i)};\Theta+\Delta\Theta)||^{2}.$ (5)
During continual domain expansion, this regularization term ensures that each previous domain query consistently generates the same parameters, with $\Theta^{*}$ denoting the hypernetwork parameters at the end of the last domain expansion,
@@ III-B Normalizing Flow for Domain Awareness
Specifically, we implement a normalizing flow using a series of affine coupling layers, a technique employed by several previous works [63, 64], denoted as $f=f^{1}\circ f^{2}\circ\ldots\circ f^{K}$. With the input split into two subsets as $[h_{a},h_{b}]=h$, each coupling layer executes an invertible transformation:
@@ III-C Deep Bayesian Uncertainty Estimation
Continual domain expansion is triggered when incoming domain datasets lack a specialized model capable of delivering satisfactory performance during training. A similar challenge can arise during inference if none of the currently expanded specialized models perform effectively. Thanks to the domain awareness provided by the normalizing flow, we can effectively detect such situations when all flow models output low likelihoods, indicating an out-of-distribution scenario. To address this, the final component of our proposed DECODE framework introduces a performance lower bound, utilizing the generalized model to ensure coverage in these scenarios.
To enhance our framework, we integrate deep Bayesian uncertainty estimation, specifically employing the concept of natural posterior networks by [28]. We start by adapting the learning objective for motion prediction. Assume that the future motion $y$ is modeled as a variable distributed according to an exponential family distribution:
@@
Together with the fact that the conjugate prior distribution $\mathbb{Q}(\eta)$ from the same exponential family can be parameterized as:
$\mathbb{Q}(\eta;\chi^{\tt{prior}},e^{\tt{prior}})=\exp{(e^{\tt{prior}}\eta^{T}\chi^{\tt{prior}}-e^{\tt{prior}}A(\eta))},$ (15)
where $\chi^{\tt{prior}}$ and $e^{\tt{prior}}$ are the prior parameters and prior evidence, respectively. The posterior distribution then becomes:
@@
with $\chi^{\tt{post}}=\frac{e^{\tt{prior}}\chi^{\tt{prior}}+\sum_{y\in D}u(y)}{e^{\tt{prior}}+|D|}$ and $e^{\tt{post}}=e^{\tt{prior}}+|D|$, where $|D|$ denotes the cardinality of the domain dataset. This results in the posterior parameters being a weighted combination of the prior parameters and the sufficient statistics of the domain data samples.
The mathematical formulations described previously lay the foundation for establishing a performance lower bound within our framework, which is achieved by using a pre-trained generalized model to predict the prior parameters. Accordingly, the decoder, as referenced in (2), is adapted to output these predicted parameters $\chi$. Given $m^{*}$ as the ID of the best specialized model as determined by (12), the posterior parameters are updated by integrating the predicted parameters from both the generalized and specialized models. The specific equations for this integration process are as follows:
@@
[…]
@@
16:          Sample parameters $\eta^{k}\sim\mathbb{Q}(\eta|\chi^{\tt{post}},e^{\tt{post}})$triangleright Alternatively compute the mean parameter $\bar{\eta}$ for deterministic sampling
@@ IV-A Motion Prediction Model Implementation
The original MTR model outputs future motions as a distribution represented by a Gaussian Mixture Model (GMM). Following 13 in the last section, the posterior distribution can be similarly modeled as a GMM, which also belongs to the exponential family. Given that a GMM combines component selection, represented as a categorical distribution, with trajectory generation for each component modeled as a multivariate Gaussian distribution, we have simplified our approach for practicality. During training, we found it sufficient to perform the Bayesian update only for the component selection part, while maintaining the original settings for trajectory generation. For the categorical distribution, the conjugate prior is formulated as a Dirichlet distribution $\eta\sim\tt{Dir}(\alpha)$, where $\alpha=e\chi$. Regarding trajectory sampling, commonly only a limited number of future trajectories are generated to represent the distribution of future motion. Thus, we combine the components of the generalized and specialized models, and employ non-maximum suppression (NMS) to select the components along with their corresponding predicted trajectories.
@@ V-A Experimental Setup
Method $\tt{minADE}_{\tt{D1,\_}}$ $\tt{minFDE}_{\tt{D1,\_}}$ $\tt{minADE}_{\tt{D2,\_}}$ $\tt{minFDE}_{\tt{D2,\_}}$ $\tt{minADE}_{\tt{D3,\_}}$ $\tt{minFDE}_{\tt{D3,\_}}$
@@
oriMTR $0.470\rightarrow 9.008\rightarrow 2.487$ $1.153\rightarrow 21.639\rightarrow 6.668$ $0.442\rightarrow 23.505$ $1.202\rightarrow 53.436$ $0.758$ $1.987$
@@
ewcMTR $0.520\rightarrow 2.345\rightarrow 1.819$ $1.224\rightarrow 6.276\rightarrow 5.025$ $0.364\rightarrow 6.429$ $0.960\rightarrow 16.827$ $0.755$ $2.030$
erMTR $\textbf{0.470}\rightarrow 1.680\rightarrow 1.313$ $\textbf{1.153}\rightarrow 4.112\rightarrow 3.146$ $\textbf{0.443}\rightarrow 0.778$ $\textbf{1.121}\rightarrow 1.751$ 0.712 1.889
DECODE $0.523\rightarrow\textbf{0.525}\rightarrow\textbf{0.578}$ $1.262\rightarrow\textbf{1.268}\rightarrow\textbf{1.319}$ $0.520\rightarrow\textbf{0.595}$ $1.262\rightarrow\textbf{1.278}$ $0.765$ $1.982$
@@
Method $\tt{AER}$ $\tt{FGT}$
($\tt{minADE}$ / $\tt{minFDE}$) ($\tt{minADE}$ / $\tt{minFDE}$)
oriMTR $6.110$ / $14.335$ $11.205$ / $26.057$
preMTR $0.959$ / $2.426$ $0.0$ / $0.0$
ewcMTR $2.038$ / $5.390$ $3.063$ / $8.240$
erMTR $0.896$ / $2.193$ $0.789$ / $1.860$
DECODE 0.584 / 1.395 0.044 / 0.030

### v2

@@ Abstract
Motion prediction is essential for autonomous vehicles to navigate complex environments and anticipate the behavior of other traffic participants. As new driving scenarios emerge, models must be continually updated without retraining from scratch. We propose DECODE, a continual learning framework that starts from a pre-trained generalized model and incrementally expands specialized models for distinct domains. Unlike existing approaches that pursue a single unified model, DECODE explicitly balances specialization and generalization through dynamic model selection. It employs a hypernetwork for parameter generation, which reduces storage costs, and utilizes a normalizing flow for real-time domain inference via likelihood estimation. Outputs from specialized and generalized models are fused using Bayesian uncertainty estimation. This integration ensures optimal performance in familiar conditions while maintaining robustness in novel scenarios. Extensive experiments show DECODE achieves a low forgetting rate of 0.044 and an average minADE of 0.584 m, outperforming prior methods and generalizing well across diverse driving domains. Furthermore, we demonstrate that DECODE can be extended beyond motion prediction to general continual learning tasks such as image classification, showcasing its broad applicability.
@@ I Introduction
Motion prediction has attracted significant attention in recent years, finding applications across a wide array of fields. It plays a critical role in the development of autonomous vehicles, enabling them to understand complex scenarios and anticipate the future behaviors of other traffic participants [34, 15]. This capability allows for more informed decision-making, which is crucial not only for enhancing roadway safety but also for improving traffic efficiency and reducing congestion. Beyond road transport, motion prediction is instrumental in areas such as robotics for navigation and interaction [23, 24], and in simulation environments where predictive models support the development and testing of various autonomous systems [4, 75]. The task of motion prediction is inherently challenging due to the complexity of interactions among target agents and the multi-modal nature of their possible future movements.
@@
Recent advancements in state-of-the-art algorithms have focused on integrating complex multi-agent interactions and map data [61, 80, 29, 63], yielding successful performances on several prominent open datasets [9, 78, 19]. However, these models require extensive datasets covering a broad spectrum of interactions involving multiple agents and road topology. Recent studies [20] have shown that without further training, models that excel on one dataset may perform poorly on others, posing a challenge for autonomous vehicles’ global deployment and operation in diverse, unforeseen scenarios. Traditionally, motion prediction development has involved training models on entire datasets—a method still prevalent in current state-of-the-art efforts. However, the nature of autonomous driving development, characterized by the continuous acquisition of new and varied driving scenarios, necessitates frequent updates through model retraining. Retraining on all data combined becomes progressively more resource-intensive, while retraining solely on new data can lead to the loss of previously acquired knowledge [73]. This situation underscores the critical need for ongoing innovation in motion prediction methodologies and the development of effective continual learning strategies to enhance vehicle safety and reliability across varied environments. Therefore, in this paper we propose a novel continual learning framework for motion prediction called Domain-awarE COntinual Domain Expansion (DECODE).
@@
Extensive experiments confirm the effectiveness and superior performance of the DECODE framework. By leveraging a hypernetwork for continual learning, we achieved a notably low forgetting rate of 0.044, surpassing several other active learning strategies, and an average minADE of just 0.584 meters. Our normalizing flow-based model identification method accurately distinguishes among different domains, providing robust evidence of its ability to align model selection with specific operational scenarios, achieving an AUROC of 0.988. We further validated our framework using both naturalistic trajectory data and synthetically generated data, demonstrating its capacity to consistently produce high-quality motion predictions across a diverse range of domains, including safety-critical scenarios, with an accuracy of 0.986. Although the framework is developed with motion prediction in mind, its core design is not task-specific, and we demonstrate its potential applicability to general continual learning problems through additional experiments.
@@ II-A Motion Prediction
Neural network-based learning approaches have significantly outperformed traditional model-based methods such as constant velocity models and Kalman filters in motion prediction [44]. These neural networks typically employ an encoder-decoder structure. During the encoding phase, techniques such as recurrent neural networks and 1D convolutions take advantage of the sequential nature of data to encode past trajectories. Recent research has further advanced motion prediction by focusing on two critical scene contexts: multi-agent interactions and local map features. Initial pioneering efforts, including Social LSTM [1] and Social GANs [30], explored the impact of multi-agent interactions on future behaviors. Meanwhile, local maps have typically been rasterized into 2D images [8, 66, 61], allowing CNNs to extract detailed scene representations from them.
@@
Subsequent advances in Graph Neural Networks (GNNs) and transformers have refined the capture of scene context information. LaneGCN [48] effectively discretizes the local map into lane nodes along lane centerlines, creating a connected lane graph that utilizes graph convolutions to extract relational information among agents and lanes—an approach echoed by several later models [77, 27]. Additionally, VectorNet [25], widely adopted in recent research [80, 29, 67], encodes both agents and map elements as polylines into a global graph. A graph neural network is then employed to encode this graph, effectively integrating the spatial relationships within the scene. Originally utilized in natural language processing, transformers have recently gained significant traction in motion prediction [49, 28, 35, 63]. Like graph neural networks, they can flexibly extract relational information among agents and map elements by leveraging the multi-head attention mechanism [68].
@@
Given the possibility of multiple future outcomes, many studies have incorporated multi-modal predictions to approximate the distribution of potential future motions of a target agent. These approaches typically generate a discrete set of future predictions, each with a corresponding confidence score to capture outcome variability. Early methods [66, 48] used multiple regression heads to produce unique modes. Generative models like GANs [30, 60] and CVAEs[61, 14] are also widely employed. More recent methods [55, 51, 7] often involve an intermediate classification phase to identify a distinct future mode, followed by regression to refine the motion. For example, TNT [80] and denseTNT [29] use agent’s target endpoints to represent future modes, GOHOME [27] predicts future motions from a heatmap, and MTR++ [63] employ a learnable intention query with k-means clustered intention points. Our DECODE framework is compatible with any motion prediction model that utilizes an encoder-decoder structure and is flexible regarding input format, allowing for the integration of state-of-the-art scene context encoding methods. Furthermore, our framework adeptly handles the mult-modality of predictions by acknowledging the potential for multiple viable future trajectories and is particularly effective with methods that incorporate an intermediate classification phase.
@@ II-B Continual Learning
Continual learning can be broadly categorized into four approaches. Parameter isolation methods, such as progressive neural networks [59], expand untouched model parameters with the arrival of new domain data, completely protecting previous knowledge at the cost of increased model complexity and parameter storage. Regulation-based methods mitigate forgetting by adding a regularization term to the loss function [2]; an example is Elastic Weight Consolidation (EWC) [38], which employs Fisher information to protect important parameters related to old knowledge during training. Knowledge distillation-based methods, such as Learning without Forgetting (LwF) [47], leverage outputs from old models to distill knowledge into the new model, preserving prior learning while integrating new information. Lastly, rehearsal-based methods include experience replay, which utilizes a selective collection from older datasets, and generative replay, which uses synthetically generated samples from previously learned models.
Recently, continual learning with pre-trained models has gained increasing attention. Traditional continual learning methods train models from scratch, which is inefficient and overlooks the prior knowledge embedded in pre-trained models. Building upon transfer learning, parameter-efficient tuning (PET) methods—such as prompt/prefix tuning [45, 46], adapter tuning [32, 56], and LoRA [33] —adapt large pre-trained models with only a small number of learnable parameters. These techniques have been extended to continual learning to mitigate catastrophic forgetting while maintaining efficiency. Representative examples include L2P [71] and DualPrompt [70], which maintain prompt pools and use query-based selection, and CODA-Prompt [65], which decomposes prompts and employs attention-based assembly for end-to-end optimization. Side-tuning [79] extends adapter-based methods but needs task identity during inference, and [26] offers a unified PET framework for continual learning.
@@
These existing approaches present several drawbacks. Experience replay struggles with selecting representative samples and necessitates external storage, which can quickly become unmanageable [54]. Generative replay’s effectiveness heavily relies on the quality of synthetic data, posing significant challenges for state-of-the-art motion prediction models that require handling complex inputs like varying agent numbers and intricate road geometry. Moreover, training one AI model using synthetically generated data from another can degrade performance, a risk highlighted by Shumailov et al. in a recent Nature study [64]. To address these issues, our DECODE framework employs hypernetworks [69], which innovatively combine parameter isolation with the regularization of loss function. Similar to parameter-efficient tuning methods, our approach begins with a pre-trained generalized model and incrementally expands a set of specialized models based on it, enabling efficient adaptation while preserving prior knowledge. This approach efficiently preserves the integrity of previous models’ parameters while integrating new knowledge, offering a scalable and effective solution to the challenges of continual learning in motion prediction.
@@ III-A Hypernetworks for Continual Domain Expansion
[…]
During continual domain expansion, this regularization term ensures that each previous domain query consistently generates the same parameters, with $\Theta^{*}$ denoting the hypernetwork parameters at the end of the last domain expansion.
@@ III-B Normalizing Flow for Domain Awareness
Specifically, we implement a normalizing flow using a series of affine coupling layers, a technique employed by several previous works [16, 36], denoted as $f=f^{1}\circ f^{2}\circ\ldots\circ f^{K}$. With the input split into two subsets as $h_{a},h_{b}]=h$, each coupling layer executes an invertible transformation:
@@ III-C Deep Bayesian Uncertainty Estimation
Continual domain expansion is triggered when incoming domain datasets lack a specialized model capable of delivering satisfactory performance during training. A similar challenge can arise during inference if none of the currently expanded specialized models perform effectively. For the task of motion prediction, this situation is particularly common. Thanks to the domain awareness provided by the normalizing flow, we can effectively detect such situations when all flow models output low likelihoods, indicating an out-of-distribution scenario. Since multi-modal motion prediction models must produce a set of future trajectories that jointly capture diverse motion intentions and sufficiently cover likely behaviors, selection mechanisms must consider both spatial diversity and the associated uncertainty of each trajectory. To this end, we propose to incorporate the uncertainty of the trajectories from the specialized models into the selection process, allowing us to identify underperforming predictions and fall back on generalist outputs when needed. The final component of our proposed DECODE framework introduces a performance lower bound, utilizing the generalized model to ensure coverage in these scenarios.
To do so, we incorporate deep Bayesian uncertainty estimation based on Natural Posterior Networks [11]. We begin by modeling the future motion $y$ using an exponential family distribution:
@@
Given that the conjugate prior distribution $\mathbb{Q}(\eta)$ belongs to the same exponential family, the posterior can be parameterized as:
@@
The updated posterior parameters are given by $\chi^{\tt{post}}=\frac{e^{\tt{prior}}\chi^{\tt{prior}}+\sum_{y\in D}T(y)}{e^{\tt{prior}}+|D|}$ and $e^{\tt{post}}=e^{\tt{prior}}+|D|$, with $\chi^{\tt{prior}}$ and $e^{\tt{prior}}$ denoting the prior parameters and prior evidence, respectively, and $|D|$ the size of the domain dataset. This formulation yields a weighted combination of the prior and the observed sufficient statistics, effectively balancing prior knowledge with new evidence.
The above formulation provides the foundation for defining a performance lower bound in our framework, achieved by using a pre-trained generalized model to predict the prior parameters. The decoder in (2) is adapted to output these predicted $\chi$ values. Given $m^{*}$ as the ID of the best specialized model from (12), the posterior parameters are then computed by combining predictions from both the generalized and specialized models. The integration is defined as:
@@
6:  $e^{(m^{*})}=\underset{m}{\tt{max}}\mathbb{P}_{h}(h;\mathcal{H}_{\omega}(q^{(m)};\Theta_{\omega}))$
@@
16:    Sample parameters $\eta^{k}\sim\mathbb{Q}(\eta|\chi^{\tt{post}},e^{\tt{post}})$ $\triangleright$ Alternatively compute the mean parameter $\bar{\eta}$ for deterministic sampling
@@ IV-A Motion Prediction Model Implementation
The original MTR model outputs future motions as a distribution represented by a Gaussian Mixture Model (GMM). Following 13 in the last section, the posterior distribution can be similarly modeled as a GMM. Given that a GMM combines component selection, represented as a categorical distribution, with trajectory generation for each component modeled as a multivariate Gaussian distribution, we have simplified our approach for practicality. During training, we found it sufficient to perform the Bayesian update only for the component selection part, while maintaining the original settings for trajectory generation. For the categorical distribution, the conjugate prior is formulated as a Dirichlet distribution $\eta\sim\tt{Dir}(\alpha)$, where $\alpha=e\chi$. Regarding trajectory sampling, commonly only a limited number of future trajectories are generated to represent the distribution of future motion. Thus, we combine the components of the generalized and specialized models, and employ non-maximum suppression (NMS) to select the components along with their corresponding predicted trajectories.
@@ V-A Experimental Setup
Method $\tt{minADE}_{\tt{D1,\_}}\downarrow$ $\tt{minFDE}_{\tt{D1,\_}}\downarrow$ $\tt{minADE}_{\tt{D2,\_}}\downarrow$ $\tt{minFDE}_{\tt{D2,\_}}\downarrow$ $\tt{minADE}_{\tt{D3,\_}}\downarrow$ $\tt{minFDE}_{\tt{D3,\_}}\downarrow$
@@
oriMTR $\textbf{0.470}\rightarrow 9.008\rightarrow 2.487$ $\textbf{1.153}\rightarrow 21.639\rightarrow 6.668$ $0.442\rightarrow 23.505$ $1.202\rightarrow 53.436$ $0.758$ $1.987$
@@
EWC $0.520\rightarrow 2.345\rightarrow 1.819$ $1.224\rightarrow 6.276\rightarrow 5.025$ $\textbf{0.364}\rightarrow 6.429$ $\textbf{0.960}\rightarrow 16.827$ $0.755$ $2.030$
MAS $0.517\rightarrow 3.330\rightarrow 1.759$ $1.240\rightarrow 8.300\rightarrow 4.436$ $0.395\rightarrow 5.120$ $1.059\rightarrow 13.18$ $0.768$ $2.030$
ER(500) $0.470\rightarrow 1.680\rightarrow 1.313$ $1.153\rightarrow 4.112\rightarrow 3.146$ $0.443\rightarrow 0.778$ $1.121\rightarrow 1.751$ $0.712$ $1.889$
DER++(500) $0.470\rightarrow 0.781\rightarrow 0.713$ $1.153\rightarrow 1.764\rightarrow 1.669$ $0.421\rightarrow 0.492$ $1.096\rightarrow 1.354$ $0.711$ $1.929$
Dualprompt $0.547\rightarrow 0.591\rightarrow 0.618$ $1.290\rightarrow 1.381\rightarrow 1.398$ $0.447\rightarrow\textbf{0.452}$ $1.217\rightarrow\textbf{1.229}$ $0.996$ $2.541$
Coda-P $0.494\rightarrow 0.710\rightarrow 0.976$ $1.173\rightarrow 1.632\rightarrow 2.263$ $0.371\rightarrow 0.899$ $1.060\rightarrow 2.488$ 0.648 1.681
DECODE $0.523\rightarrow\textbf{0.525}\rightarrow\textbf{0.578}$ $1.262\rightarrow\textbf{1.268}\rightarrow\textbf{1.319}$ $0.520\rightarrow 0.595$ $1.262\rightarrow 1.278$ $0.765$ $1.982$
@@
[…]